40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:03:02 2016, sisyphus, Metaphysics, SUS, Broken wire Unknown.jpeg
    Reply  Wed Jul 27 16:56:47 2016, rana, Metaphysics, SUS, Broken wire 
    Reply  Thu Jul 28 08:42:36 2016, Steve, Update, SUS, wire clamp e12342.jpg
    Reply  Fri Jul 29 15:17:23 2016, Steve, Update, SUS, wire and clamping test 43micr242grs.jpg
       Reply  Fri Jul 29 20:12:54 2016, gautam, Update, SUS, third time - no luck 
          Reply  Mon Aug 1 14:09:20 2016, Steve, Update, SUS, third time - no luck 
             Reply  Tue Aug 9 16:19:27 2016, Steve, Update, SUS, torque driver for wire clamp 
                Reply  Wed Aug 10 15:34:24 2016, Steve, Update, SUS, 6 in-lbs torque driver for wire clamp screw 
          Reply  Wed Aug 3 00:15:39 2016, gautam, Update, SUS, ETMX suspended 
             Reply  Wed Aug 3 09:26:54 2016, Lydia, Update, SUS, ETMX suspended: photos 8x
             Reply  Thu Aug 4 03:14:39 2016, gautam, Update, SUS, ETMX suspended 
                Reply  Fri Aug 5 02:17:10 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 8x
                   Reply  Fri Aug 5 04:43:09 2016, Koji, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 
                   Reply  Fri Aug 5 18:21:28 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff GLUED 11x
                      Reply  Sun Aug 7 14:53:39 2016, ericq , Update, SUS, ETMX Standoff gluing was successful ETMX_resonances.pdfETMX_SUSspectra.zip
                         Reply  Tue Aug 9 00:44:43 2016, gautam , Update, SUS, ETMY patch-up 
                            Reply  Tue Aug 9 15:27:57 2016, gautam , Update, SUS, ETMY patch-up IMG_2996.JPGIMG_2997.JPG
                               Reply  Wed Aug 10 03:08:03 2016, gautam , Update, SUS, ETMY patch-up 
                                  Reply  Wed Aug 10 23:45:03 2016, gautam , Update, SUS, ETMY re-suspended 7x
                                     Reply  Thu Aug 11 00:20:41 2016, Koji, Update, SUS, ETMY re-suspended 
                                        Reply  Thu Aug 11 11:56:40 2016, gautam, Update, SUS, ETMY re-suspended pitchBalancingWithBias.PDF
                                           Reply  Mon Aug 15 18:28:15 2016, gautam, Update, SUS, Air-bake preparation IMG_6816.JPG
                                              Reply  Tue Aug 16 08:04:49 2016, Steve, Update, SUS, ETMY had Aluminum bridge AlBridge.jpg
                                              Reply  Tue Aug 16 21:54:27 2016, gautam, Update, SUS, Air-bake - IN PROGRESS IMG_3005.JPG
                                                 Reply  Wed Aug 17 08:47:43 2016, ericq, Update, SUS, Air-bake finished 
                                                    Reply  Wed Aug 17 23:00:57 2016, gautam, Update, SUS, Air-bake of towers IMG_3006.JPG
                                                       Reply  Thu Aug 18 08:17:16 2016, Steve, Update, SUS, wire clamping preparation 800P.jpg
                                                       Reply  Thu Aug 18 14:14:20 2016, gautam, Update, SUS, Air-bake of towers - finished 
                                                          Reply  Fri Aug 19 22:51:12 2016, gautam, Update, SUS, ETMs first-contacted 
                                                             Reply  Sun Aug 21 16:23:05 2016, gautam, Update, SUS, ETMX suspended IMG_3018.JPG
                                                                Reply  Mon Aug 22 13:06:11 2016, gautam, Update, SUS, ETMY suspended IMG_3019.JPG
                                                                   Reply  Mon Aug 22 18:04:24 2016, gautam, Update, SUS, ETMY OSEMs inserted 
                                                                      Reply  Tue Aug 23 19:35:38 2016, gautam, Update, SUS, ETMY back in IFO 10x
                                                                         Reply  Thu Aug 25 21:11:43 2016, gautam, Update, SUS, ETMY back in IFO ETMY_BounceSpectra_25Aug2016.pdf
                                                                            Reply  Fri Aug 26 11:35:44 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_BounceSpectra_26Aug2016.pdf41.png19.png
                                                                               Reply  Fri Aug 26 14:05:38 2016, Steve, Update, SUS, ETMY UL sensor problematic  ETMY-ULhistory.png
                                                                               Reply  Fri Aug 26 14:09:46 2016, ericq, Update, SUS, ETMY UL sensor problematic  
                                                                                  Reply  Sun Aug 28 21:40:11 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_UL_problematic.png
                                                                                     Reply  Mon Aug 29 09:37:05 2016, gautam, Update, SUS, ETMY UL sensor problematic  satelliteBox.png
                                                                                        Reply  Mon Aug 29 14:42:06 2016, gautam, Update, SUS, ETMY Satellite box diagnostics IMG_6840.JPGIMG_6841.JPG
                                                                            Reply  Fri Aug 26 17:48:14 2016, gautam, Update, SUS, bounce mode coupling reduction ETMY_BounceSpectra_26Aug2016_1.pdf
                                                                               Reply  Mon Aug 29 21:53:14 2016, Lydia, Update, SUS, ETMY bounce mode coupling IMG_3051.JPG42.png
                                                                                  Reply  Tue Aug 30 20:26:36 2016, gautam, Update, SUS, ITMY installed IMG_3052.JPGSRM_sensor_level_comp.pngETMY_BounceSpectra_30Aug2016.pdf
                                                                                     Reply  Wed Aug 31 18:07:43 2016, Johannes, Update, SUS, ITMY free swinging 
                                                                                        Reply  Wed Aug 31 22:09:18 2016, gautam, Update, SUS, Y arm locked to Green  
                                                                                           Reply  Thu Sep 1 08:30:24 2016, Steve, Update, SUS, ITMY_UL is sick ITMY_SENSOR_UL.pngspareSatAmp.jpg
                                                                                              Reply  Tue Sep 6 20:30:56 2016, rana, Update, SUS, ITMY_UL is sick 
                                                                                                 Reply  Tue Sep 13 11:00:59 2016, Lydia, Update, SUS, ETMY UL glitch returned 
                                                                                           Reply  Thu Sep 1 19:59:22 2016, Johannes, Update, SUS, IR mode flashes in Y arm 
                                                                                              Reply  Fri Sep 2 21:09:08 2016, gautam, Update, SUS, Y arm locked in air 03.png56.png
                                                                                                 Reply  Tue Sep 6 20:45:14 2016, gautam, Update, SUS, X arm test masses back in chamber IMG_3137.JPGIMG_3143.JPGIMG_3142.JPGIMG_3148.JPG
                                                                                                    Reply  Thu Sep 8 22:12:36 2016, gautam, Update, SUS, X arm in place, locked to green, IR flashes visible 
                                                                                                       Reply  Fri Sep 9 17:50:02 2016, gautam, Update, SUS, Heavy doors on BS-PRM, ETMY chambers 
                                                                                                          Reply  Sat Sep 10 14:26:01 2016, rana, Update, General, PRM SRM alignment Osamu.JPG
                                                                                                             Reply  Mon Sep 12 17:15:22 2016, Johannes, Update, General, PRM SRM alignment IMG_3273.JPGIMG_3275.JPG
                                                                                        Reply  Thu Sep 1 19:18:14 2016, Johannes, Update, SUS, ITMX and ETMX preemptive table leveling 8x
                                                                                           Reply  Fri Sep 2 21:15:53 2016, gautam, Update, SUS, ITMX and ETMX preemptive table leveling 
                                                                                     Reply  Thu Sep 1 15:43:16 2016, rana, Update, SUS, OSEM adjustments 
                                                                                        Reply  Mon Sep 5 19:57:24 2016, Osamu, Update, SUS, OSEM adjustments 
Message ID: 12466     Entry time: Fri Sep 2 21:09:08 2016     In reply to: 12465     Reply to this: 12474
Author: gautam 
Type: Update 
Category: SUS 
Subject: Y arm locked in air 

[johannes, lydia, gautam]

  • The Y arm has been locked to IR in air using POY11 as an error signal yes
  • We had been seeing flashes in the arm since yesterday, but were unable to lock
  • Today we re-did the alignment procedure much the same way as yesterday
    • It is useful to put in the slide-on irides onto the suspension tower for this sort of alignment
    • We were a bit more systematic in aligning back-reflections to overlap each other today
    • It is useful to stick the IR card just in front of the iris, and align the tip tilts by looking at the scatter on the camera. At least for Yaw, this works pretty well, probably a more reliable reference than contorting oneself inside the vacuum chamber to see if we are well aligned or not.
  • Two fixes that made locking possible today:
    • The POY error signal had a large DC offset. I zeroed the offset and adjusted the demod phase to make the error signal 0 when the IMC was unlocked
    • I replaced the 50-50 beam splitter that was dividing the transmitted light between the QPD and Thorlabs PDs with a 2" Y1 CVI mirror - this meant that the flashes we had with the arm roughly aligned went from < 0.1 to a healthier 0.25, which allowed easier locking
  • The POY whitening gain was unchanged from when we locked the Y arm in air just after venting and before taking the doors off
  • The mode is barely visible on ITMY face, although I guess this is to be expected given we are at low power
  • Lydia then tuned the arm alignment more finely such that the transmission is now ~0.65 (See Attachemnt 2 for slider values)
  • From values from normal (pre-vent) IFO operation, I would have expected us to get a transmission of about 1 assuming 100mW going into the IFO from the IMC - and so with the BS switched out for an HR mirror, a transmission of ~2. What we get is about 1/3 of this value. Perhaps the IMC isn't so well aligned, but it is hard to imagine we have only 30mW going into the IFO. Or perhaps the input pointing is sub optimal (I did not run ASS, perhaps I should have)

GV EDIT Sep 5: These numbers do make sense if the ND filter that was on the Transmon QPD had ND = 0.6 (there are two at the end, one labelled ND 0.6 and the other labelled ND10 though the latter label looks like some custom label so I don't really trust that value), even though only one was on, unfortunately I don't remember which. So, for 10% of input power with a factor of 8 increase because the ND filter is removed and also that the 50% BS has been replaced with a HR mirror, we expect a transmission level of ~0.6 (compared to the normalized value under normal IFO operation) which is close to what we see...

  • The UL coil problems continue to plague us but we were able to lock the arm regardless

In any case, I think we can work on putting in the X arm now and work on recovering that. 

To do for the Y-arm (now that the F.C. is off, we should try and do this in as few chamber openings as possible):

  • Fix problematic ITMY UL coil
  • Rotation of all 5 ITMY OSEM coils for B-R peak reduction in sensor outputs
  • Adjustment of axial position of all OSEM coils on ITMY and ETMY to better center the PD outputs to half their saturation value, given that the pitch and yaw biases to the optics have changed since this was last done
  • Insertion of new baffles - try and center the IR and green beams as best as possible on these so that they serve as an alignment reference in the future

Then we need to do all of this for the X arm as well. The PRM LR coil is still giving no output - I will try moving the bias sliders around to see if this is a stuck magnet situation, but perhaps it is not. Since Eric's 3-satellite-box-monte did not yield any positive results, we have to consider the possibility that the LED or PD themselves are damaged. If so, I don't see any workaround without opening up the BS-PRM chamber, but if we can avoid this, we should. Perhaps when ITMX is open we can use the camera with the IR filter removed to see if all the OSEM LEDs are functional through the beam tube.


We are also piping POY11 error to the DAFI model and can hear it in the control room.


Rana suggested reviving the MC autolocker - I've made some changes to the low power MC autolocker scripts and they've been working the few times I tried today evening, but let's see how it does over the weekend. I've also changed the Y axis of the StripTool on the wall to better reflect the low-power range..

Attachment 1: 03.png  28 kB  | Hide | Hide all
03.png
Attachment 2: 56.png  20 kB  Uploaded Fri Sep 2 22:28:13 2016  | Hide | Hide all
56.png
ELOG V3.1.3-